OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

OpenStudy (anonymous):

If x = -4, the factor has to be (x+4) and if x = -1, the factor has to be (x+1). So the two factors are (x+4)(x+1), hence the equation has to be what?

OpenStudy (anonymous):

A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5